OpenStudy (goformit100):

Find all triples of natural numbers (a,b,c) such that a, b and c are in geometric progession, and a+b+c=111.

mathslover (mathslover):

If a,b , c are in GP then : b = ar c = br \(b^2 = ac\) --(1)

OpenStudy (goformit100):

ok

mathslover (mathslover):

We have to find, all the solns for a, b and c such that a, b and c are in GP , right?

OpenStudy (goformit100):

yes..

mathslover (mathslover):

Any ideas?

mathslover (mathslover):

@UnkleRhaukus @.Sam. @Callisto @mathstudent55 - any of u can help?

OpenStudy (goformit100):

No idea but can we solve it using pascal triangle ?

mathslover (mathslover):

It is for determining Binomial coefficients... what is the use of binomial theorem related topics in this question?

OpenStudy (goformit100):

ok

mathslover (mathslover):

I got something : \(\bf{a + c = 111 - \sqrt{ac} \\ (a+c)^2 = (111-\sqrt{ac})^2 }\)

OpenStudy (goformit100):

Ya ok

mathslover (mathslover):

\(\cfrac{111}{2} \ge b^{\frac{3}{2}} \)

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

@UnkleRhaukus

mathslover (mathslover):

@hartnn

mathslover (mathslover):

Correction : \(\cfrac{a+b+c}{3} \ge( {\sqrt{abc}})^{\frac{1}{3}}\)

OpenStudy (goformit100):

ok

mathslover (mathslover):

37+37+37 = 111 and they all r in GP too :P :)

mathslover (mathslover):

a=b=c=37

OpenStudy (goformit100):

Ok Thank you

ganeshie8 (ganeshie8):

1, 10, 100

mathslover (mathslover):

^

mathslover (mathslover):

Many numbers are there

mathslover (mathslover):

\(a(1+r+r^2) = 3 *37\) a = 3 , then 1 + r + r^2 = 37 similarly factorize 111 and do that

ganeshie8 (ganeshie8):

yea and this is a olympiad problem it wud be tough

mathslover (mathslover):

Yep.

ganeshie8 (ganeshie8):

thats a brilliant logic :)

OpenStudy (goformit100):

Thank you for the help :)

ganeshie8 (ganeshie8):

111, -111, 111

mathslover (mathslover):

@ganeshie8 , -111 is not a natural number. So, 111,-111,111 is not possible. We only have 2 sets of solns : (37,37,37) and (1,10,100) \(a+ar + ar^2 = 111\) \(a(1 + r + r^2) = 111\) Case 1 : \(a(1+r+r^2) = 1 \times 111\) a = 1 or \(1 + r + r^2 = 1 \) and a = 111 or \(1+ r + r^2 = 111\) \(1 + r + r^2 \) can not be equal to 1. So, a = 1 and \(1 + r + r^2 = 111\) \(1 + r + r^2 = 111\) \(r^2 + r -110 = 0\) This gives, r = 10 (r can not be negative) If r = 10, then we have : a = 1 , b = ar = 10, c = ar^2 = 100 \(\bf{(1,10,100)}\) Case 2 : \(a(1+r+r^2) = 3 \times 37\) Either a =3 or \(1 + r+r^2 =3\) and a = 37 or \(1 + r + r^2 = 37\) \(1 + r +r^2 \ne 37\) as , r can not be zero. Therefore, a = 37 and \(1 + r + r^2 =3\) \(r^2 + r + 1 = 3\) Which gives, r = 1 (r can not be negative) Therefore , we get : a = 37, b = ar = 37 , c = ar^2 = 37 \(\bf{(37,37,37)}\)